OpenStudy (anonymous):

A bag cotains 3 blue marbles and 5 red marbles samuel chooses a marble at random then without replacing the marble he chooses another marble at random what is the probability that he chose 2 red marbles

OpenStudy (anonymous):

There are 8 marbles in the bag. The chance of picking a red ball is 5 out of the 8 - 5/8 On the second go, you have one less red ball. So, the chance of picking a red ball is 4 out of the remaining 7 - 4/7 In simple probability, and=* multiply or=+ add So its the first probability AND the second probability. Did that help?

OpenStudy (anonymous):

yes it did but i have another one for you can you help

OpenStudy (anonymous):

charles's mp3 player has 100 songs on it of those 40 are hip hop 30 are rock and 30 are rap the mp3 player is set to random and once a song has been played it cannot be played again what is the probability that the first song he listens to is a rap song and the second song is a hip hop song

OpenStudy (anonymous):

So, it's like the marble problem. What's the probability that the first song is Rap? It's the number of Rap songs over the total number of songs. Then find the probability for the second part, using the number of Hip Hop songs out of the remaining songs. Then we either need to Add or Multiply the two together. What do you think?

OpenStudy (anonymous):

ok i am confused sorry

OpenStudy (anonymous):

So the first part of the question is: Pick out a Rap song, from all the songs. There are 20 Rap songs and 100 songs in total. 20 Rap songs, out of 100 song. So like the marbles (where it was 5 red marbles out of 8 - 5/8) here is 20 rap songs out of 100 songs - 20/100 (which is 1/5 when we cancel down). Now we are left with 99 songs (as that song can be played twice. The second part asks us to pick out a Hip Hip songs. Whats the probability of picking out a Hip Hop song from the songs that are left?

OpenStudy (anonymous):

so would it be 99/100 or like 1/4

OpenStudy (anonymous):

There are 99 songs left (100-1) There are 20 Hip Hop songs. The probability is 20/99 (how many of what we want divided by how many are left in total). Does that make sense?

OpenStudy (anonymous):

well i divided it by 20 and 99 then also as a fraction and got for both of them 0.20202020202

OpenStudy (anonymous):

I see what you're trying to do but you're making a bit more complicated. Its 20 Rap songs out of the 100 songs, then 20 Hip Hop songs out of the remaining 99 songs. (20/100)AND(20/99) = (20/100)*(20/99)=0.04040404... = 0.04

OpenStudy (anonymous):

k so a spiner has 11 equal sized spaces with 7 spaces red and 4 blue a number cube is numbered 1 to 6 the girl spins the spinner and rolls the number cube what is the probability that the spinner lands on red and the umber cube shows an odd number

OpenStudy (anonymous):

a 7/22 b 2/11 c 7/66 or d 2/33